OpenStudy (anonymous):

WILL GIVE MEDALS find a quadratic equation with the roots -1+4i -1-4i

OpenStudy (anonymous):

-1+4i -1-4i

hartnn (hartnn):

hi :) did you know that if 'a' is a root, then (x-a) is a factor of that function.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

just really lost on how to do this

hartnn (hartnn):

now you know :) so ,when -1+4i is a root, (x-[-1+4i]) is a factor. right ? what about the -1-4i as the root ?

OpenStudy (anonymous):

(x+[1-4i])???

hartnn (hartnn):

nopes, we just put \((x-root)\) so maybe try again once ?

OpenStudy (anonymous):

(x-[-1-4i])?

hartnn (hartnn):

correct!

hartnn (hartnn):

so your equation is just (x- [-1+4i]) (x- [-1-4i]) = 0 now you just need to foil the left side, can you ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x^2+2x+17=0 be another way of writing this?

hartnn (hartnn):

x^2+2x+17=0 is correct!

OpenStudy (anonymous):

really? lol wow i feel so dumb XD

OpenStudy (anonymous):

so would that be the answer?

hartnn (hartnn):

yes, it is correct final answer :) good work!

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcome ^_^